OpenStudy (anonymous):

Solve -3(7x-6)-x=26-2(11x=4)

OpenStudy (anonymous):

show work

OpenStudy (anonymous):

confirm the sign should not be =4

OpenStudy (anonymous):

excuse my mistake its +4

OpenStudy (anonymous):

-3(7x -6) -x = 26 -2(11x + 4) -21x + 18 -x = 26 - 22x - 8 -22x + 18 = -22x + 18 -22x +22x = 18 -18 0 = 0 unless i read that wrong not correct

OpenStudy (anonymous):

let me know - I have looked at this twice and i get 0 = 0

OpenStudy (anonymous):

im looking at it and thats all i could come up with also

OpenStudy (anonymous):

so basically this is an identity solution. Well we both are correct :)

OpenStudy (anonymous):

so technically x can be anything :)

OpenStudy (anonymous):

thank you much :)